OpenStudy (anonymous):

Make x the subject: x + 3 = 2a + bx

OpenStudy (anonymous):

\[x-bx = 2a-3\]\[x(1-b) = 2a-3\]\[x = \frac{2a-3}{1-b}\]

OpenStudy (anonymous):

mr. idk anything?